OpenStudy (anonymous):

Simplify (7/(w+3))+((w+7)/(w^2-9))+(6/(w-3))

OpenStudy (anonymous):

\[\frac{7}{w+3}+\frac{w+7}{w^2-9}+\frac{6}{w-3}\]

OpenStudy (anonymous):

\[\frac{7}{w+3}+\frac{w+7}{(w+3)(w-3)}+\frac{6}{w-3}\]

OpenStudy (anonymous):

Is that as far as you can go?

OpenStudy (anonymous):

\[\frac{7(w-3)+(w+7)+6(w+3)}{(w+3)(w-3)}\]

OpenStudy (anonymous):

alright that's what i thought. i wasn't sure if i could make any more simplifications